# Hyperelliptic four-manifolds defined by vector-colorings of simple polytopes

### v2

@@ 4.7. Geometric hyperelliptic manifolds $N(P,\Lambda)$
In dimension $n=4$ among all products of Euclidean, spherical and hyperbolic geometries (including these geometries) the geometries $\mathbb{S}^{4}$, $\mathbb{S}^{3}\times\mathbb{R}$, $\mathbb{S}^{2}\times\mathbb{S}^{2}$, $\mathbb{S}^{2}\times\mathbb{R}^{2}$, $\mathbb{S}^{2}\times\mathbb{L}^{2}$, and $\mathbb{L}^{2}\times\mathbb{L}^{2}$ admit hyperelliptic manifolds build by Construction 4.45, and the geometries $\mathbb{R}^{4}$, $\mathbb{L}^{4}$, $\mathbb{L}^{3}\times\mathbb{R}$, $\mathbb{L}^{2}\times\mathbb{R}^{2}$ do not admit.
@@
The cases of $\mathbb{S}^{4}$, $\mathbb{S}^{3}\times\mathbb{R}$, $\mathbb{S}^{2}\times\mathbb{S}^{2}$, $\mathbb{S}^{2}\times\mathbb{R}^{2}$, $\mathbb{S}^{2}\times\mathbb{L}^{2}$, $\mathbb{L}^{2}\times\mathbb{L}^{2}$ are covered by Propositions 4.58, 4.61, Corollaries 4.68, 4.74, and Example 4.75.
The cases of $\mathbb{R}^{4}$, $\mathbb{L}^{4}$, $\mathbb{L}^{3}\times\mathbb{R}$, $\mathbb{L}^{2}\times\mathbb{R}^{2}$ are covered by Corollary 4.49, and Propositions 4.56, 4.57. ∎
@@ 5. Acknowledgements
The authors are grateful to V.M. Buchstaber for his permanent attention, to A.D. Mednykh for motivating discussions, in particular for the suggestion to consider geometries arising on hyperelliptic $4$-manifolds $N(P,\Lambda)$ and the corresponding branch sets, and to D.V. Gugnin for useful discussions.

### v3

@@ 4.7. Geometric hyperelliptic manifolds $N(P,\Lambda)$
In dimension $n=4$ among all products of Euclidean, spherical and hyperbolic geometries (including these geometries) the geometries $\mathbb{S}^{4}$, $\mathbb{S}^{3}\times\mathbb{R}$, $\mathbb{S}^{2}\times\mathbb{S}^{2}$, $\mathbb{S}^{2}\times\mathbb{R}^{2}$, $\mathbb{S}^{2}\times\mathbb{L}^{2}$, and $\mathbb{L}^{2}\times\mathbb{L}^{2}$ admit hyperelliptic manifolds build by Construction 4.46, and the geometries $\mathbb{R}^{4}$, $\mathbb{L}^{4}$, $\mathbb{L}^{3}\times\mathbb{R}$, $\mathbb{L}^{2}\times\mathbb{R}^{2}$ do not admit.
@@
The cases of $\mathbb{S}^{4}$, $\mathbb{S}^{3}\times\mathbb{R}$, $\mathbb{S}^{2}\times\mathbb{S}^{2}$, $\mathbb{S}^{2}\times\mathbb{R}^{2}$, $\mathbb{S}^{2}\times\mathbb{L}^{2}$, $\mathbb{L}^{2}\times\mathbb{L}^{2}$ are covered by Propositions 4.59, 4.62, Corollaries 4.69, 4.75, and Example 4.76.
The cases of $\mathbb{R}^{4}$, $\mathbb{L}^{4}$, $\mathbb{L}^{3}\times\mathbb{R}$, $\mathbb{L}^{2}\times\mathbb{R}^{2}$ are covered by Corollary 4.50, and Propositions 4.57, 4.58. ∎
@@ 5. Acknowledgements
The study has been funded within the framework of the HSE University Basic Research Program.
The author is grateful to V. M. Buchstaber for his permanent attention, to A. D. Mednykh for motivating discussions, in particular for the suggestion to consider geometries arising on hyperelliptic $4$-manifolds $N(P,\Lambda)$ and the corresponding branch sets, and to D. V. Gugnin, A. Yu. Vesnin, and V. Yu. Gorchakov for useful discussions.